### v1

# MaLA-500: Massive Language Adaptation of Large Language Models

Abstract
Large language models have advanced the state of the art in natural language processing. However, their predominant design for English or a limited set of languages creates a substantial gap in their effectiveness for low-resource languages. To bridge this gap, we introduce MaLA-500, a novel large language model designed to cover an extensive range of 534 languages. To train MaLA-500, we employ vocabulary extension and continued pretraining on LLaMA 2 with Glot500-c. Our experiments on SIB-200 show that MaLA-500 achieves state-of-the-art in-context learning results. We release MaLA-500 at https://huggingface.co/MaLA-LM.
**footnotetext: Equal contribution.
1 Introduction
Large Language Models (LLMs), e.g., LLaMA (Touvron et al., 2023a, b), Mistral (Jiang et al., 2023, 2024), and ChatGPT, [^1] https://openai.com/blog/chatgpt have shown remarkable performance in natural language understanding and generation. Follow-up studies (Bang et al., 2023; Lai et al., 2023; Ahuja et al., 2023a, b) observe that these English-centric LLMs, such as LLaMA with mainly English as the training data, are capable of handling some high-resource non-English languages, benefiting from the inclusion of non-English language data during pretraining. However, their applicability to low-resource languages is still limited due to data scarcity and model capability in terms of model size.
Recently, several generative multilingual LLMs, such as XGLM (Lin et al., 2021), mGPT (Shliazhko et al., 2022), and BLOOM (Scao et al., 2022), have emerged. Notably, the current language coverage for these LLMs is limited to up to 60 languages, highlighting the remaining need for further work on massively multilingual LLMs.
ImaniGooghari et al. (2023) have achieved a significant milestone in the realm of massive language adaptation by extending the language coverage of a small-scale multilingual language model, XLM-R (Conneau et al., 2020) - an auto-encoding model with 278M parameters, from 100 languages to an impressive number of 534 languages, and introducing an extended model, Glot500-m with 395M parameters. ImaniGooghari et al. (2023) introduce the Glot500-c corpora spanning 534 languages from 47 language families, and then apply vocabulary extension and continued pretraining to create Glot500-m. The introduction of Glot500-c significantly mitigates the challenge of data scarcity for low-resource languages. Moreover, the adaptation method is more favorable than training from scratch, as it requires fewer computational resources and emits a smaller carbon footprint. This success serves as a strong motivation for our exploration into the massive language adaptation of LLMs.
This work aims to extend the capabilities of LLMs to encompass a wider range of languages. Existing works on language adaptation of pretrained models provide extended coverage across a wide linguistic spectrum but are limited to relatively small model sizes - mostly at the hundred million scale. Our study pushes the boundaries by exploring language adaptation techniques for LLMs with model parameters scaling up to 10B. Our investigation delves into generative LLMs with a substantial increase in model parameters and their in-context learning capabilities that enhance contextual and linguistic relevance across a diverse range of languages. We address the challenges of adapting LLMs to low-resource languages, such as data sparsity, domain-specific vocabulary, and linguistic diversity. We study continued pretraining of open LLM (i.e., LLaMA 2 (Touvron et al., 2023b)), vocabulary extension, and adaptation techniques (i.e., LoRA low-rank reparameterization (Hu et al., 2022)). We train and release MaLA-500 that covers more than 500 languages in various domains. We evaluate MaLA-500 on SIB-200 and the results show that MaLA-500 outperforms existing open LLMs of close or slightly larger model size. This work broadens the accessibility of LLMs, making them valuable for a more diverse set of language-specific use cases, especially for low-resource ones, and addressing the equality issue by removing language barriers for speakers of many languages, especially those underrepresented languages covered by existing LLMs.
2 Massive Language Adaptation
The principle of massive language adaption of large language models accommodates the utilization of a massively multilingual corpus (Section 2.1), the strong base LLM (Section 2.2), and the technique for effective language adaption (Sections 2.3 and 2.4).
2.1 Data
We use Glot500-c (ImaniGooghari et al., 2023) covering 534 languages [^2] We define languages using the ISO 639-3 code combined with the corresponding written script. For example, “eng_Latn” represents English written in the Latin script. as the training data of MaLA-500. See §A for the list of languages with their data amounts. We sample languages from the imbalanced dataset according to a multinomial distribution, with $\alpha=0.3$ for vocabulary extension and continued pretraining.
2.2 Model
We choose LLaMA 2 (Touvron et al., 2023b) to start continual training. LLaMA series models (Touvron et al., 2023a), with model weights released publicly, have gained popularity in the research community. Despite being English-centric compared to their multilingual counterparts, they have shown remarkable capacity for multiple languages (Zhao et al., 2024). We choose the latest LLaMA 2, trained on 2 trillion tokens, as our base model to benefit from its outstanding language capacity. Our study chooses the 7B model with 32 transformer layers.
2.3 Vocabulary Extension
The original LLaMA 2’s 32,000 tokenizer covers English and a small fraction of other European languages using Latin or Cyrillic scripts. To enhance its capability and encoding efficiency for a broader range of languages, we extend the vocabulary with Glot500-c. Specifically, we initially train a multilingual tokenizer with SentencePiece (Kudo and Richardson, 2018) on the sampled Glot500-c with a vocabulary of 250,000. Subsequently, we merge the trained tokenizer with the original LLaMA 2 tokenizer by taking the union of their vocabularies. As a result, we obtain the MaLA-500’s tokenizer with a vocabulary size of 260,164. After vocabulary extension and resizing the embedding layer, the model size becomes 8.6B.
We measure the impact of vocabulary extension on the development set of Glot500-c by analyzing the reduction in segmentation length for each language. The results indicate that the effect of vocabulary extension varies, ranging from 8% (English, eng_Latn) to 88% (Oriya, ori_Orya). Unsurprisingly, vocabulary extension has a larger effect on languages written in non-Latin scripts than on those in the Latin script. However, for some low-resource languages written in the Latin script, e.g., Kabiyè (kbp_Latn) and Vietnamese (vie_Latn), the segmentation length is shortened by around 50%.
2.4 Continued Pretraining
We employ continued pretraining for language adaptation with low-rank adaptation (LoRA, Hu et al., 2022) to enable parameter-efficient training, given the limitation of our computing resources. LoRA injects trainable rank decomposition matrices, which approximate the large weight matrices with a lower rank, to the pretrained model weights. It significantly reduces the computational complexity and thus saves the training cost. We continually train the casual language model to update the rank-decomposition matrices while freezing the model weights of pretrained models, allowing the continually trained language model to learn from new data in new languages without completely losing its previous language capacity. Continual training of large language models requires substantial computational resources. We adopt efficient distributed training setups on supercomputers to make the training process feasible.
2.5 Training
Hardware and Software
We train our model on the computing cluster with the theoretical peak performance of 2 petaflops on GPU nodes. We deploy distributed training on 24 Nvidia Ampere A100 GPUs. As for software, we utilize the Huggingface Transformers (Wolf et al., 2020), PEFT (Parameter-Efficient Fine-Tuning) [^3] https://huggingface.co/docs/peft/index, and DeepSpeed (Rasley et al., 2020). We use the ZeRO redundancy optimizer (Rajbhandari et al., 2020) and maximize the batch size that fits the memory of each GPU. We employ mixed-precision training using the bfloat16 format.
Hyperparameters
The learning rate is set at 2e-4. A weight decay of 0.01 is applied to penalize large weights and mitigate overfitting. The trainable LoRA module targets all linear layers, excluding the language model head. In our setting, the final model has 10B parameters in total, in which 2B parameters are trainable. The LoRA module is incorporated with a rank of 8, an alpha value of 32, and a dropout rate of 0.05, contributing to the model’s adaptability and regularization during training. The context window is 2k. We maximize the batch size to fit the memory, making a global batch size of 384. The model undergoes three training epochs. Checkpoints are saved every 500 steps, and we employ early stopping to select the checkpoint that exhibits the most favorable average performance on downstream tasks. In this model release, we only consider the topic classification task.
Environmental Impacts
We train our model on CSC’s carbon-neutral data center, with all electricity generated with renewable hydropower, [^4] https://www.csc.fi/sustainable-development and the waste heat is utilized in district heating to further reduce CO2 footprint. [^5] https://www.csc.fi/sustainability-1
3 Evaluation
We evaluate the few-shot learning capability of MaLA-500 and compare it with other LLMs on a topic classification dataset of SIB-200 (Adelani et al., 2023). The classification task involves seven classes, namely science/technology, travel, politics, sports, health, entertainment, and geography. Our evaluation spans a diverse set of 177 languages, obtained by intersecting the language sets of SIB-200 and Glot500-c. In this evaluation, we compute the accuracy of 99 samples for each language, with shots randomly sampled from the respective training sets.
The evaluated LLM receives a structured prompt, which is the concatenation of few-shot examples and the sample intended for prediction. The format for both a few-shot example and the sample to predict is defined as follows:
Topic Classification: science/technology, travel, politics, sports, health, entertainment, geography.
The label of [sent] is [label]
where [sent] is the sentence for topic classification, and [label] is the ground truth. [label] is included if the sample serves as a few-shot example or omitted if it is the sample to predict. The constructed prompt is then used as input to the LLM. Subsequently, the evaluated LLM is prompted to estimate the probability of the label over the label set based on the provided prompt. This evaluation process is implemented using the lm-evaluation-harness. [^6] https://github.com/EleutherAI/lm-evaluation-harness
Table 1: 3-shot in-context learning macro-average accuracy (%) on SIB-200 of different LLMs. mGPT has no model with around 7B parameters, so we choose a larger one with 13B parameters.
Model Accuracy
LLaMA 2-7B 41.94
mGPT-13B 39.96
BLOOM-7B1 44.22
XGLM-7.5B 47.53
MaLA-500 54.10
[Uncaptioned image] Figure 1: Detailed results of 3-shot in-context learning on SIB-200. X-axis: the number of languages in different accuracy ranges (%).
Table 1: 3-shot in-context learning macro-average accuracy (%) on SIB-200 of different LLMs. mGPT has no model with around 7B parameters, so we choose a larger one with 13B parameters.
3.1 Comparison across LLMs
We compare MaLA-500 with LLaMA 2-7B, mGPT-13B, BLOOM-7B1, and XGLM-7.5B, focusing on the 3-shot in-context learning macro-average accuracy across languages. As shown in Table. 1, MaLA-500 outperforms previous LLMs significantly. Notably, MaLA-500 exhibits a remarkable 12.16% improvement compared to our base model, LLaMA 2, highlighting MaLA-500’s substantial contribution to enhancing the multilingual capacity of LLMs.
Fig. 1 provides a detailed performance analysis across languages. Notably, MaLA-500 surpasses previous LLMs in the sense that no language obtains random results (0-20%) and the number of languages with very low accuracy (20-40%) is significantly reduced. Moreover, MaLA-500 excels in the number of languages achieving 60% accuracy or higher (65 altogether).
In our comprehensive analysis of contributing factors, a significant Pearson correlation of 0.53 is observed between the effect of vocabulary extension and the performance gain. This observation holds true for languages with both non-Latin scripts, such as Kannada (kan_Knda), Malayalam (mal_Mlym), and Tigrinya (tir_Ethi), as well as Latin scripts, such as Igbo (ibo_Latn) and Yoruba (yor_Latn). Furthermore, we note that the corpus size of a language exhibits a weak correlation of 0.13 with its performance gain. In contrast, the corpus size of the language family to which a language belongs demonstrates a moderate correlation of 0.40.
3.2 Effect of Number of Shots
Fig. 2 illustrates the relationship between accuracy and the number of in-context examples (i.e., shots). As the number of in-context shots increases, there is a corresponding rise in accuracy. Notably, with just 1-shot, accuracy exhibits randomness at 26.08%, indicating 1-shot provides limited information for task learning. The transition from 1 shot to 2 shots/3 shots results in a significant improvement, with performances boosted by 20.97% and 28.02%, respectively. This highlights the effectiveness of increasing the number of shots. MaLA-500 achieves its peak performance at approximately 60% accuracy with 6-10 in-context shots. This may be attributed to the multi-class nature of the SIB-200 dataset, necessitating more shots for learning intricate input-label mappings.
In Fig. 3, a more nuanced portrayal of results aligns with the observations made in Fig. 2. In the realm of 1-shot in-context learning, approximately 40 languages exhibit erratic results. As the number of shots increases, there is a reduction in the number of languages achieving low accuracy (20-40%), coupled with a growing cohort achieving high accuracy (60-80%). Notably, in the case of 10-shot in-context learning, 101 (57%) languages surpass the threshold of 60% accuracy.
Further examination into individual language trends reveals that some low-resource languages require more shots to achieve better performance (e.g., pes_Arab for Persian) or even exhibit poor performance with 10 shots (e.g., dzo_Tibt for Dzongkha and ayr_Latn for Central Aymara). In contrast, high-resource languages, such as fra_Latn for French, demonstrate impressive performance even with fewer shots, and increasing the number of shots results in only marginal improvement.
Figure 2: In-context learning macro-average accuracy (%) on SIB-200 with different number of shots using MaLA-500.
Figure 3: Detailed results of in-context learning on SIB-200 using MaLA-500. X-axis: the number of languages in different accuracy ranges (%). Y-axis: number of shots.
4 Related Work
4.1 Multilingual Language Models
Language model development has endeavored to broaden the scope of pertaining languages to address multilingual scenarios. Pretrained multilingual models have been able to accommodate up to a hundred or more languages. Noteworthy examples include mBERT Devlin et al. (2019), which supports 104 languages, XLM-R (Conneau et al., 2020) covering 100 languages, mBART (Liu et al., 2020) designed for 25 languages, mT5 (Xue et al., 2021) spanning 101 languages, XGLM (Lin et al., 2021) across 30 languages, GPT-3 covering 118 languages (93% English), mGPT (Shliazhko et al., 2022) accommodating 60 languages, and BLOOM (Scao et al., 2022) supporting 46 languages and 13 programming languages.
Surprisingly, two recent multilingual language models have surpassed the conventional limit by supporting more than 400 languages. Glot500-m (ImaniGooghari et al., 2023) spans 511 languages through vocabulary extension and continued training based on XLM-R. SERENGETI (Adebara et al., 2022) goes even further by supporting 517 African languages and language varieties, written in five different scripts, employing models inspired by both ELECTRA (Clark et al., 2020) and XLM-R. MADLAD (Kudugunta et al., 2023) covers 419 languages and trains an 8B language model from scratch with an adapted UL2 objective (Tay et al., 2022). Our work is concurrent with the MADLAD400 language model. We distinguish it from: 1) language coverage. Our work covered more than 500 languages, a number comparable to that of encoder-only models and surpassing MADLAD-400 by an additional 100 languages. 2) training methods. We consider continual training to benefit from the learned knowledge of the original models. 3) model architecture. We adopt an open model architecture, i.e., LLaMA, while MADLAD uses UL2.
4.2 Language Adaptation
Before the advent of LLMs, diverse approaches are employed to adapt small-scale multilingual language models to new languages. These methods include using adapters (Pfeiffer et al., 2020; Üstün et al., 2020; Pfeiffer et al., 2020; Nguyen et al., 2021; Faisal and Anastasopoulos, 2022; Yong et al., 2022), vocabulary extension and substitution (Chau et al., 2020; Wang et al., 2020; Müller et al., 2020, 2021; Pfeiffer et al., 2021; Chen et al., 2023b; Downey et al., 2023), leveraging monolingual corpora (Ebrahimi and Kann, 2021), and utilizing bilingual lexicons (Wang et al., 2022).
While language models have been scaled up significantly, their coverage is limited to a specific set of languages. To address this constraint, various methods have been proposed to expand the applicability of these Large Language Models (LLMs) across a broader range of languages, catering to both general-purpose tasks and specific applications like machine translation. These methods also involve vocabulary extension (Cui et al., 2023), continued pretraining and instruction-tuning (Yong et al., 2022; Cui et al., 2023; Chen et al., 2023a), and parallel corpora exploitation (Cahyawijaya et al., 2023; Yang et al., 2023; Zhu et al., 2023; Xu et al., 2023). Despite these efforts, massive language adaptation of LLMs for general-purpose tasks across diverse languages remains an area yet to be thoroughly explored.
5 Conclusion
We present a pioneering effort in massive language adaptation on LLMs, focusing on extending LLaMA 7B to our model, MaLA-500. This adaptation involves vocabulary extension and continued pretraining with LoRA. Our approach leads to MaLA-500 achieving state-of-the-art in-context learning capabilities, as demonstrated on the task of SIB-200. We release the model weights publicly to facilitate future research. This work marks a substantial advancement in applying LLMs to a diverse range of languages.
Limitations
Data
We do not collect more high-resource data and instead use only Glot500-c, which does not put full effort on collecting data for high-resource languages. In addition, the flores200-based SIB-200 evaluation set is included in the training data since Glot500-c includes flores200, but the classification labels are not provided.
Model Size
Our study limits the maximum model size to 10B parameters. In our future work, we plan to expand to a larger model.
Ethical Implications
LLMs have been known to exhibit biases present in their training data. When extending LLMs to low-resource languages, there is a risk of propagating biases from high-resource languages to underrepresented ones. Careful attention must be paid to mitigate bias and ensure fairness in data collection and model training. The paper aims to make LLMs more accessible for underrepresented languages. Still, there is a risk of creating a digital language divide if certain communities are left out due to limited technological access. Future work would address biases by conducting bias audits on the training data, debiasing the models during generation, and continuously monitoring model outputs.
Acknowledgements
This work was funded by the European Research Council (grants #740516 and #758969), EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), and the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546]. The authors wish to acknowledge CSC – IT Center for Science, Finland, for generous computational resources. Shaoxiong Ji and Peiqin Lin acknowledge travel support from ELISE (GA no 951847).
Appendix A Languages
The list of languages of Glot500-c used to train MaLA-500 with the number of available sentences and language family information for each language is available in Tables 2,  3 and 4.
Lang |Sent| Family Lang |Sent| Family Lang |Sent| Family
hbs_Latn 63411156 indo1319 hin_Deva 7046700 indo1319 ton_Latn 1216118 aust1307
mal_Mlym 48098273 drav1251 kor_Hang 6468444 kore1284 tah_Latn 1190747 aust1307
aze_Latn 46300705 ory_Orya 6266475 indo1319 lat_Latn 1179913 indo1319
guj_Gujr 45738685 indo1319 urd_Arab 6009594 indo1319 srn_Latn 1172349 indo1319
ben_Beng 43514870 indo1319 swa_Latn 5989369 ewe_Latn 1161605 atla1278
kan_Knda 41836495 drav1251 sqi_Latn 5526836 indo1319 bem_Latn 1111969 atla1278
tel_Telu 41580525 drav1251 bel_Cyrl 5319675 indo1319 efi_Latn 1082621 atla1278
mlt_Latn 40654838 afro1255 afr_Latn 5157787 indo1319 bis_Latn 1070170 indo1319
fra_Latn 39197581 indo1319 nno_Latn 4899103 indo1319 orm_Latn 1067699
spa_Latn 37286756 indo1319 tat_Cyrl 4708088 turk1311 haw_Latn 1062491 aust1307
eng_Latn 36122761 indo1319 ast_Latn 4683554 indo1319 hmo_Latn 1033636 pidg1258
fil_Latn 33493255 aust1307 mon_Cyrl 4616960 mong1349 kat_Geor 1004297 kart1248
nob_Latn 32869205 indo1319 hbs_Cyrl 4598073 indo1319 pag_Latn 983637 aust1307
rus_Cyrl 31787973 indo1319 hau_Latn 4368483 afro1255 loz_Latn 964418 atla1278
deu_Latn 31015993 indo1319 sna_Latn 4019596 atla1278 fry_Latn 957422 indo1319
tur_Latn 29184662 turk1311 msa_Latn 3929084 mya_Mymr 945180 sino1245
pan_Guru 29052537 indo1319 som_Latn 3916769 afro1255 nds_Latn 944715 indo1319
mar_Deva 28748897 indo1319 srp_Cyrl 3864091 indo1319 run_Latn 943828 atla1278
por_Latn 27824391 indo1319 mlg_Latn 3715802 pnb_Arab 899895 indo1319
nld_Latn 25061426 indo1319 zul_Latn 3580113 atla1278 rar_Latn 894515 aust1307
ara_Arab 24524122 arz_Arab 3488224 afro1255 fij_Latn 887134 aust1307
zho_Hani 24143786 nya_Latn 3409030 atla1278 wls_Latn 882167 aust1307
ita_Latn 23539857 indo1319 tam_Taml 3388255 drav1251 ckb_Arab 874441 indo1319
ind_Latn 23018106 aust1307 hat_Latn 3226932 indo1319 ven_Latn 860249 atla1278
ell_Grek 22033282 indo1319 uzb_Latn 3223485 turk1311 zsm_Latn 859947 aust1307
bul_Cyrl 21823004 indo1319 sot_Latn 3205510 atla1278 chv_Cyrl 859863 turk1311
swe_Latn 20725883 indo1319 uzb_Cyrl 3029947 turk1311 lua_Latn 854359 atla1278
ces_Latn 20376340 indo1319 cos_Latn 3015055 indo1319 que_Latn 838486
isl_Latn 19547941 indo1319 als_Latn 2954874 indo1319 sag_Latn 771048 atla1278
pol_Latn 19339945 indo1319 amh_Ethi 2862985 afro1255 guw_Latn 767918 atla1278
ron_Latn 19190217 indo1319 sun_Latn 2586011 aust1307 bre_Latn 748954 indo1319
dan_Latn 19174573 indo1319 war_Latn 2584810 aust1307 toi_Latn 745385 atla1278
hun_Latn 18800025 ural1272 div_Thaa 2418687 indo1319 pus_Arab 731992 indo1319
tgk_Cyrl 18659517 indo1319 yor_Latn 2392359 atla1278 che_Cyrl 728201 nakh1245
srp_Latn 18371769 indo1319 fao_Latn 2365271 indo1319 pis_Latn 714783 indo1319
fas_Arab 18277593 uzn_Cyrl 2293672 turk1311 kon_Latn 685194
ceb_Latn 18149215 aust1307 smo_Latn 2290439 aust1307 oss_Cyrl 683517 indo1319
heb_Hebr 18128962 afro1255 bak_Cyrl 2264196 turk1311 hyw_Armn 679819 indo1319
hrv_Latn 17882932 indo1319 ilo_Latn 2106531 aust1307 iso_Latn 658789 atla1278
glg_Latn 17852274 indo1319 tso_Latn 2100708 atla1278 nan_Latn 656389 sino1245
fin_Latn 16730388 ural1272 mri_Latn 2046850 aust1307 lub_Latn 654390 atla1278
slv_Latn 15719210 indo1319 hmn_Latn 1903898 lim_Latn 652078 indo1319
vie_Latn 15697827 aust1305 asm_Beng 1882353 indo1319 tuk_Latn 649411 turk1311
mkd_Cyrl 14717004 indo1319 hil_Latn 1798875 aust1307 tir_Ethi 649117 afro1255
slk_Latn 14633631 indo1319 nso_Latn 1619354 atla1278 tgk_Latn 636541 indo1319
nor_Latn 14576191 indo1319 ibo_Latn 1543820 atla1278 yua_Latn 610052 maya1287
est_Latn 13600579 kin_Latn 1521612 atla1278 min_Latn 609065 aust1307
ltz_Latn 12997242 indo1319 hye_Armn 1463123 indo1319 lue_Latn 599429 atla1278
eus_Latn 12775959 oci_Latn 1449128 indo1319 khm_Khmr 590429 aust1305
lit_Latn 12479626 indo1319 lin_Latn 1408460 atla1278 tum_Latn 589857 atla1278
kaz_Cyrl 12378727 turk1311 tpi_Latn 1401844 indo1319 tll_Latn 586530 atla1278
lav_Latn 12143980 indo1319 twi_Latn 1400979 atla1278 ekk_Latn 582595 ural1272
bos_Latn 11014744 indo1319 kir_Cyrl 1397566 turk1311 lug_Latn 566948 atla1278
epo_Latn 8737198 arti1236 pap_Latn 1360138 indo1319 niu_Latn 566715 aust1307
cat_Latn 8648271 indo1319 nep_Deva 1317291 indo1319 tzo_Latn 540262 maya1287
tha_Thai 7735209 taik1256 azj_Latn 1315834 turk1311 mah_Latn 534614 aust1307
ukr_Cyrl 7462046 indo1319 bcl_Latn 1284493 aust1307 tvl_Latn 521556 aust1307
tgl_Latn 7411064 aust1307 xho_Latn 1262364 atla1278 jav_Latn 516833 aust1307
sin_Sinh 7293178 indo1319 cym_Latn 1244783 indo1319 vec_Latn 514240 indo1319
gle_Latn 7225513 indo1319 gaa_Latn 1222307 atla1278 jpn_Jpan 510722 japo1237
Table 2: List of languages of Glot500-c (Part I).
Lang |Sent| Family Lang |Sent| Family Lang |Sent| Family
lus_Latn 509250 sino1245 kmb_Latn 296269 atla1278 ncx_Latn 162558 utoa1244
crs_Latn 508755 indo1319 zai_Latn 277632 otom1299 qug_Latn 162500 quec1387
kqn_Latn 507913 atla1278 gym_Latn 274512 chib1249 rmn_Latn 162069 indo1319
ndo_Latn 496613 atla1278 bod_Tibt 273489 sino1245 cjk_Latn 160645 atla1278
snd_Arab 488730 indo1319 nde_Latn 269931 atla1278 arb_Arab 159884 afro1255
yue_Hani 484700 sino1245 fon_Latn 268566 atla1278 kea_Latn 158047 indo1319
tiv_Latn 483064 atla1278 ber_Latn 264426 mck_Latn 157521 atla1278
kua_Latn 473535 atla1278 nbl_Latn 259158 atla1278 arn_Latn 155882 arau1255
kwy_Latn 473274 atla1278 kmr_Latn 256677 indo1319 pdt_Latn 155485 indo1319
hin_Latn 466175 indo1319 guc_Latn 249044 araw1281 her_Latn 154827 atla1278
iku_Cans 465011 mam_Latn 248348 maya1287 gla_Latn 152563 indo1319
kal_Latn 462430 eski1264 nia_Latn 247406 aust1307 kmr_Cyrl 151728 indo1319
tdt_Latn 459818 aust1307 nyn_Latn 241992 atla1278 mwl_Latn 150054 indo1319
gsw_Latn 449240 indo1319 cab_Latn 240101 araw1281 nav_Latn 147702 atha1245
mfe_Latn 447435 indo1319 top_Latn 239232 toto1251 ksw_Mymr 147674 sino1245
swc_Latn 446378 atla1278 tog_Latn 231969 atla1278 mxv_Latn 147591 otom1299
mon_Latn 437950 mong1349 mco_Latn 231209 mixe1284 hif_Latn 147261 indo1319
mos_Latn 437666 atla1278 tzh_Latn 230706 maya1287 wol_Latn 146992 atla1278
kik_Latn 437228 atla1278 pms_Latn 227748 indo1319 sme_Latn 146803 ural1272
cnh_Latn 436667 sino1245 wuu_Hani 224088 sino1245 gom_Latn 143937 indo1319
gil_Latn 434529 aust1307 plt_Latn 220413 aust1307 bum_Latn 141673 atla1278
pon_Latn 434522 aust1307 yid_Hebr 220214 indo1319 mgr_Latn 138953 atla1278
umb_Latn 431589 atla1278 ada_Latn 219427 atla1278 ahk_Latn 135068 sino1245
lvs_Latn 422952 indo1319 iba_Latn 213615 aust1307 kur_Arab 134160 indo1319
sco_Latn 411591 indo1319 kek_Latn 209932 maya1287 bas_Latn 133436 atla1278
ori_Orya 410827 koo_Latn 209375 atla1278 bin_Latn 133256 atla1278
arg_Latn 410683 indo1319 sop_Latn 206501 atla1278 tsz_Latn 133251 tara1323
kur_Latn 407169 indo1319 kac_Latn 205542 sino1245 sid_Latn 130406 afro1255
dhv_Latn 405711 aust1307 qvi_Latn 205447 quec1387 diq_Latn 128908 indo1319
luo_Latn 398974 nilo1247 cak_Latn 204472 maya1287 srd_Latn 127064
lun_Latn 395764 atla1278 kbp_Latn 202877 atla1278 tcf_Latn 126050 otom1299
nzi_Latn 394247 atla1278 ctu_Latn 201662 maya1287 bzj_Latn 124958 indo1319
gug_Latn 392227 tupi1275 kri_Latn 201087 indo1319 udm_Cyrl 121705 ural1272
bar_Latn 387070 indo1319 mau_Latn 199134 otom1299 cce_Latn 120636 atla1278
bci_Latn 384059 atla1278 scn_Latn 199068 indo1319 meu_Latn 120273 aust1307
chk_Latn 380596 aust1307 tyv_Cyrl 198649 turk1311 chw_Latn 119751 atla1278
roh_Latn 377067 indo1319 ina_Latn 197315 arti1236 cbk_Latn 118789 indo1319
aym_Latn 373329 ayma1253 btx_Latn 193701 aust1307 ibg_Latn 118733 aust1307
yap_Latn 358929 aust1307 nch_Latn 193129 utoa1244 bhw_Latn 117381 aust1307
ssw_Latn 356561 atla1278 ncj_Latn 192962 utoa1244 ngu_Latn 116851 utoa1244
quz_Latn 354781 quec1387 pau_Latn 190529 aust1307 nyy_Latn 115914 atla1278
sah_Cyrl 352697 turk1311 toj_Latn 189651 maya1287 szl_Latn 112496 indo1319
tsn_Latn 350954 atla1278 pcm_Latn 187594 indo1319 ish_Latn 111814 atla1278
lmo_Latn 348135 indo1319 dyu_Latn 186367 mand1469 naq_Latn 109747 khoe1240
ido_Latn 331239 arti1236 kss_Latn 185868 atla1278 toh_Latn 107583 atla1278
abk_Cyrl 321578 abkh1242 afb_Arab 183694 afro1255 ttj_Latn 106925 atla1278
zne_Latn 318871 atla1278 urh_Latn 182214 atla1278 nse_Latn 105189 atla1278
quy_Latn 311040 quec1387 quc_Latn 181559 maya1287 hsb_Latn 104802 indo1319
kam_Latn 310659 atla1278 new_Deva 181427 sino1245 ami_Latn 104559 aust1307
bbc_Latn 310420 aust1307 yao_Latn 179965 atla1278 alz_Latn 104392 nilo1247
vol_Latn 310399 arti1236 ngl_Latn 178498 atla1278 apc_Arab 102392 afro1255
wal_Latn 309873 gong1255 nyu_Latn 177483 atla1278 vls_Latn 101900 indo1319
uig_Arab 307302 turk1311 kab_Latn 176015 afro1255 mhr_Cyrl 100474 ural1272
vmw_Latn 306899 atla1278 tuk_Cyrl 175769 turk1311 djk_Latn 99234 indo1319
kwn_Latn 305362 atla1278 xmf_Geor 174994 kart1248 wes_Latn 98492 indo1319
pam_Latn 303737 aust1307 ndc_Latn 174305 atla1278 gkn_Latn 97041 atla1278
seh_Latn 300243 atla1278 san_Deva 165616 indo1319 grc_Grek 96986 indo1319
tsc_Latn 298442 atla1278 nba_Latn 163485 atla1278 hbo_Hebr 96484 afro1255
nyk_Latn 297976 atla1278 bpy_Beng 162838 indo1319 swh_Latn 95776 atla1278
Table 3: List of languages of Glot500-c (Part II).
Lang |Sent| Family Lang |Sent| Family Lang |Sent| Family
alt_Cyrl 95148 turk1311 mny_Latn 50581 atla1278 csy_Latn 34126 sino1245
rmn_Grek 94533 indo1319 gkp_Latn 50549 mand1469 azb_Arab 33758 turk1311
miq_Latn 94343 misu1242 kat_Latn 50424 kart1248 csb_Latn 33743 indo1319
kaa_Cyrl 88815 turk1311 bjn_Latn 49068 aust1307 tpm_Latn 33517 atla1278
kos_Latn 88603 aust1307 acr_Latn 48886 maya1287 quw_Latn 33449 quec1387
grn_Latn 87568 dtp_Latn 48468 aust1307 rmy_Cyrl 33351 indo1319
lhu_Latn 87255 sino1245 lam_Latn 46853 atla1278 ixl_Latn 33289 maya1287
lzh_Hani 86035 sino1245 bik_Latn 46561 mbb_Latn 33240 aust1307
ajp_Arab 83297 afro1255 poh_Latn 46454 maya1287 pfl_Latn 33148 indo1319
cmn_Hani 80745 sino1245 phm_Latn 45862 atla1278 pcd_Latn 32867 indo1319
gcf_Latn 80737 indo1319 hrx_Latn 45716 indo1319 tlh_Latn 32863 arti1236
rmn_Cyrl 79925 indo1319 quh_Latn 45566 quec1387 suz_Deva 32811 sino1245
kjh_Cyrl 79262 turk1311 hyw_Cyrl 45379 indo1319 gcr_Latn 32676 indo1319
rng_Latn 78177 atla1278 rue_Cyrl 45369 indo1319 jbo_Latn 32619 arti1236
mgh_Latn 78117 atla1278 eml_Latn 44630 indo1319 tbz_Latn 32264 atla1278
xmv_Latn 77896 aust1307 acm_Arab 44505 afro1255 bam_Latn 32150 mand1469
ige_Latn 77114 atla1278 tob_Latn 44473 guai1249 prk_Latn 32085 aust1305
rmy_Latn 76991 indo1319 ach_Latn 43974 nilo1247 jam_Latn 32048 indo1319
srm_Latn 76884 indo1319 vep_Latn 43076 ural1272 twx_Latn 32028 atla1278
bak_Latn 76809 turk1311 npi_Deva 43072 indo1319 nmf_Latn 31997 sino1245
gur_Latn 76151 atla1278 tok_Latn 42820 arti1236 caq_Latn 31903 aust1305
idu_Latn 75106 atla1278 sgs_Latn 42467 indo1319 rop_Latn 31889 indo1319
yom_Latn 74818 atla1278 lij_Latn 42447 indo1319 tca_Latn 31852 ticu1244
tdx_Latn 74430 aust1307 myv_Cyrl 42147 ural1272 yan_Latn 31775 misu1242
mzn_Arab 73719 indo1319 tih_Latn 41873 aust1307 xav_Latn 31765 nucl1710
cfm_Latn 70227 sino1245 tat_Latn 41640 turk1311 bih_Deva 31658
zpa_Latn 69237 otom1299 lfn_Latn 41632 arti1236 cuk_Latn 31612 chib1249
kbd_Cyrl 67914 abkh1242 cgg_Latn 41196 atla1278 kjb_Latn 31471 maya1287
lao_Laoo 66966 taik1256 ful_Latn 41188 atla1278 hne_Deva 31465 indo1319
nap_Latn 65826 indo1319 gor_Latn 41174 aust1307 wbm_Latn 31394 aust1305
qub_Latn 64973 quec1387 ile_Latn 40984 arti1236 zlm_Latn 31345 aust1307
oke_Latn 64508 atla1278 ium_Latn 40683 hmon1336 tui_Latn 31161 atla1278
ote_Latn 64224 otom1299 teo_Latn 40203 nilo1247 ifb_Latn 30980 aust1307
bsb_Latn 63634 aust1307 kia_Latn 40035 atla1278 izz_Latn 30894 atla1278
ogo_Latn 61901 atla1278 crh_Cyrl 39985 turk1311 rug_Latn 30857 aust1307
abn_Latn 61830 atla1278 crh_Latn 39896 turk1311 aka_Latn 30704 atla1278
ldi_Latn 61827 atla1278 enm_Latn 39809 indo1319 pxm_Latn 30698 book1242
ayr_Latn 61570 ayma1253 sat_Olck 39614 aust1305 kmm_Latn 30671 sino1245
gom_Deva 61140 indo1319 mad_Latn 38993 aust1307 mcn_Latn 30666 afro1255
bba_Latn 61123 atla1278 cac_Latn 38812 maya1287 ifa_Latn 30621 aust1307
aln_Latn 60989 indo1319 hnj_Latn 38611 hmon1336 dln_Latn 30620 sino1245
leh_Latn 59944 atla1278 ksh_Latn 38130 indo1319 ext_Latn 30605 indo1319
ban_Latn 59805 aust1307 ikk_Latn 38071 atla1278 ksd_Latn 30550 aust1307
ace_Latn 59333 aust1307 sba_Latn 38040 cent2225 mzh_Latn 30517 mata1289
pes_Arab 57511 indo1319 zom_Latn 37013 sino1245 llb_Latn 30480 atla1278
skg_Latn 57228 aust1307 bqc_Latn 36881 mand1469 hra_Latn 30472 sino1245
ary_Arab 56933 afro1255 bim_Latn 36835 atla1278 mwm_Latn 30432 cent2225
hus_Latn 56176 maya1287 mdy_Ethi 36370 gong1255 krc_Cyrl 30353 turk1311
glv_Latn 55641 indo1319 bts_Latn 36216 aust1307 tuc_Latn 30349 aust1307
fat_Latn 55609 atla1278 gya_Latn 35902 atla1278 mrw_Latn 30304 aust1307
frr_Latn 55254 indo1319 ajg_Latn 35631 atla1278 pls_Latn 30136 otom1299
mwn_Latn 54805 atla1278 agw_Latn 35585 aust1307 rap_Latn 30102 aust1307
mai_Deva 54687 indo1319 kom_Cyrl 35249 ural1272 fur_Latn 30052 indo1319
dua_Latn 53392 atla1278 knv_Latn 35196 kaa_Latn 30031 turk1311
dzo_Tibt 52732 sino1245 giz_Latn 35040 afro1255 prs_Arab 26823 indo1319
ctd_Latn 52135 sino1245 hui_Latn 34926 nucl1709 san_Latn 25742 indo1319
nnb_Latn 52041 atla1278 kpg_Latn 34900 aust1307 som_Arab 14199 afro1255
sxn_Latn 51749 aust1307 zea_Latn 34426 indo1319 uig_Latn 9637 turk1311
mps_Latn 50645 tebe1251 aoj_Latn 34349 nucl1708 hau_Arab 9593 afro1255
Table 4: List of languages of Glot500-c (Part III).
Appendix B Detailed Results
Detailed results of few-shot learning evaluation are shown in Tables 5 and  6.
Lang LLaMA 2 7B mGPT 13B BLOOM 7B1 XGLM 7.5B MaLA-500
1-shot 2-shot 3-shot 4-shot 5-shot 6-shot 7-shot 8-shot 9-shot 10-shot
ace_Latn 49.49 45.45 50.51 54.55 34.34 50.51 54.55 59.60 55.56 58.59 62.63 57.58 67.68 47.47
acm_Arab 45.45 51.52 65.66 64.65 34.34 50.51 53.54 52.53 55.56 56.57 57.58 57.58 55.56 76.77
afr_Latn 53.54 47.47 54.55 58.59 34.34 54.55 66.67 66.67 65.66 76.77 71.72 67.68 75.76 48.48
ajp_Arab 51.52 52.53 67.68 64.65 33.33 49.49 49.49 49.49 47.47 48.48 47.47 51.52 56.57 67.68
als_Latn 50.51 44.44 46.46 57.58 21.21 57.58 70.71 62.63 63.64 69.70 69.70 67.68 71.72 75.76
amh_Ethi 18.18 16.16 13.13 20.20 19.19 38.38 43.43 39.39 43.43 45.45 45.45 39.39 44.44 49.49
apc_Arab 46.46 55.56 69.70 64.65 25.25 50.51 48.48 47.47 45.45 55.56 51.52 52.53 53.54 55.56
arb_Arab 46.46 47.47 68.69 61.62 31.31 52.53 49.49 54.55 57.58 56.57 59.60 58.59 55.56 65.66
ary_Arab 42.42 42.42 67.68 61.62 24.24 46.46 47.47 45.45 44.44 49.49 45.45 46.46 50.51 74.75
arz_Arab 47.47 54.55 75.76 64.65 34.34 50.51 48.48 51.52 53.54 60.61 55.56 51.52 58.59 35.35
asm_Beng 32.32 47.47 63.64 54.55 22.22 41.41 48.48 47.47 47.47 47.47 49.49 46.46 52.53 64.65
ast_Latn 60.61 46.46 61.62 64.65 28.28 57.58 67.68 67.68 65.66 70.71 70.71 67.68 76.77 71.72
ayr_Latn 35.35 32.32 32.32 34.34 23.23 30.30 38.38 43.43 41.41 47.47 47.47 46.46 44.44 40.40
azb_Arab 40.40 37.37 37.37 41.41 27.27 38.38 38.38 39.39 45.45 51.52 40.40 43.43 45.45 70.71
azj_Latn 43.43 49.49 33.33 44.44 24.24 51.52 59.60 62.63 62.63 71.72 69.70 67.68 69.7 60.61
bak_Cyrl 35.35 54.55 34.34 44.44 34.34 49.49 61.62 63.64 63.64 70.71 64.65 63.64 65.66 72.73
bam_Latn 23.23 29.29 22.22 31.31 17.17 29.29 31.31 38.38 33.33 44.44 38.38 37.37 39.39 68.69
ban_Latn 57.58 52.53 56.57 60.61 25.25 43.43 59.60 62.63 59.60 61.62 63.64 59.6 61.62 64.65
bel_Cyrl 35.35 44.44 29.29 45.45 35.35 52.53 64.65 63.64 65.66 66.67 64.65 70.71 69.7 66.67
bem_Latn 29.29 31.31 35.35 38.38 23.23 33.33 42.42 44.44 48.48 46.46 44.44 49.49 47.47 55.56
ben_Beng 41.41 53.54 60.61 64.65 28.28 48.48 54.55 60.61 60.61 61.62 60.61 55.56 59.6 70.71
bjn_Latn 53.54 56.57 59.60 64.65 24.24 53.54 62.63 67.68 62.63 67.68 68.69 63.64 65.66 72.73
bod_Tibt 13.13 13.13 11.11 18.18 22.22 30.30 32.32 38.38 32.32 38.38 34.34 34.34 35.35 70.71
bos_Latn 65.66 54.55 46.46 55.56 27.27 59.60 69.70 64.65 64.65 70.71 75.76 68.69 75.76 72.73
bul_Cyrl 57.58 49.49 44.44 65.66 43.43 64.65 64.65 62.63 68.69 70.71 69.70 66.67 70.71 59.60
cat_Latn 61.62 52.53 57.58 66.67 29.29 57.58 68.69 61.62 65.66 68.69 69.70 66.67 72.73 66.67
ceb_Latn 61.62 48.48 48.48 49.49 29.29 61.62 69.70 73.74 66.67 80.81 74.75 72.73 74.75 60.61
ces_Latn 65.66 52.53 44.44 47.47 32.32 54.55 63.64 63.64 69.70 69.70 71.72 65.66 74.75 73.74
cjk_Latn 35.35 20.20 28.28 29.29 14.14 26.26 29.29 34.34 26.26 40.40 32.32 32.32 42.42 63.64
ckb_Arab 25.25 19.19 20.20 20.20 32.32 45.45 54.55 45.45 47.47 51.52 52.53 56.57 54.55 57.58
cmn_Hani 44.44 54.55 68.69 64.65 40.40 51.52 66.67 64.65 67.68 68.69 69.70 65.66 70.71 49.49
crh_Latn 42.42 43.43 42.42 50.51 25.25 55.56 53.54 59.60 52.53 63.64 57.58 52.53 58.59 77.78
cym_Latn 39.39 24.24 29.29 35.35 22.22 46.46 59.60 58.59 60.61 67.68 65.66 62.63 69.7 68.69
dan_Latn 60.61 53.54 53.54 61.62 31.31 52.53 65.66 65.66 67.68 74.75 74.75 71.72 73.74 34.34
deu_Latn 59.60 50.51 59.60 65.66 40.40 61.62 71.72 70.71 71.72 77.78 73.74 70.71 78.79 61.62
dyu_Latn 24.24 29.29 31.31 33.33 15.15 26.26 30.30 30.30 27.27 33.33 33.33 32.32 35.35 50.51
dzo_Tibt 16.16 12.12 15.15 17.17 19.19 32.32 28.28 28.28 28.28 29.29 31.31 35.35 32.32 72.73
ell_Grek 54.55 49.49 32.32 70.71 39.39 57.58 63.64 67.68 67.68 71.72 67.68 68.69 70.71 53.54
eng_Latn 66.67 48.48 66.67 71.72 44.44 68.69 72.73 71.72 73.74 75.76 76.77 76.77 77.78 76.77
epo_Latn 51.52 43.43 53.54 55.56 31.31 60.61 65.66 74.75 66.67 73.74 67.68 68.69 74.75 55.56
est_Latn 44.44 42.42 35.35 61.62 23.23 55.56 60.61 61.62 64.65 67.68 66.67 60.61 66.67 39.39
eus_Latn 45.45 46.46 58.59 61.62 22.22 49.49 45.45 57.58 54.55 58.59 56.57 53.54 57.58 59.60
ewe_Latn 23.23 26.26 25.25 38.38 21.21 36.36 37.37 38.38 31.31 37.37 42.42 39.39 37.37 56.57
fao_Latn 44.44 40.40 41.41 39.39 25.25 46.46 63.64 58.59 65.66 66.67 63.64 62.63 68.69 70.71
fij_Latn 30.30 30.30 33.33 42.42 17.17 34.34 41.41 45.45 46.46 50.51 46.46 45.45 48.48 39.39
fin_Latn 61.62 48.48 38.38 66.67 28.28 56.57 67.68 62.63 68.69 74.75 72.73 70.71 73.74 56.57
fon_Latn 22.22 25.25 26.26 30.30 20.20 31.31 32.32 28.28 30.30 34.34 31.31 34.34 34.34 75.76
fra_Latn 63.64 53.54 65.66 67.68 39.39 60.61 70.71 64.65 67.68 74.75 72.73 70.71 71.72 55.56
ful_Latn 35.35 31.31 34.34 44.44 18.18 34.34 44.44 41.41 45.45 49.49 56.57 46.46 58.59 59.60
fur_Latn 55.56 49.49 52.53 58.59 19.19 50.51 57.58 54.55 55.56 65.66 61.62 62.63 58.59 35.35
gla_Latn 30.30 27.27 23.23 34.34 15.15 31.31 40.40 36.36 34.34 36.36 39.39 39.39 43.43 69.70
gle_Latn 41.41 30.30 21.21 28.28 22.22 41.41 51.52 56.57 56.57 58.59 63.64 53.54 61.62 39.39
glg_Latn 56.57 45.45 63.64 65.66 31.31 54.55 69.70 64.65 72.73 75.76 78.79 74.75 81.82 41.41
grn_Latn 46.46 46.46 51.52 56.57 29.29 54.55 61.62 63.64 57.58 59.60 63.64 65.66 63.64 62.63
guj_Gujr 19.19 11.11 63.64 19.19 17.17 40.40 50.51 44.44 51.52 55.56 51.52 50.51 53.54 49.49
hat_Latn 45.45 37.37 42.42 64.65 27.27 52.53 58.59 62.63 63.64 64.65 70.71 60.61 75.76 45.45
hau_Latn 26.26 29.29 27.27 38.38 19.19 49.49 54.55 55.56 54.55 54.55 55.56 53.54 55.56 40.40
heb_Hebr 42.42 44.44 21.21 26.26 20.20 36.36 35.35 40.40 35.35 31.31 40.40 40.4 35.35 63.64
hin_Deva 51.52 51.52 63.64 70.71 24.24 42.42 53.54 58.59 57.58 61.62 60.61 57.58 56.57 65.66
hne_Deva 44.44 48.48 66.67 64.65 28.28 49.49 52.53 60.61 59.60 59.60 59.60 59.6 55.56 56.57
hrv_Latn 66.67 56.57 52.53 51.52 30.30 56.57 73.74 67.68 71.72 69.70 76.77 69.7 76.77 74.75
hun_Latn 65.66 51.52 36.36 43.43 33.33 45.45 57.58 59.60 56.57 69.70 61.62 59.6 65.66 63.64
hye_Armn 29.29 13.13 18.18 17.17 26.26 52.53 59.60 60.61 67.68 70.71 66.67 67.68 71.72 71.72
ibo_Latn 20.20 27.27 42.42 28.28 31.31 50.51 57.58 55.56 48.48 61.62 58.59 51.52 53.54 66.67
ilo_Latn 47.47 49.49 51.52 53.54 19.19 51.52 62.63 65.66 60.61 68.69 72.73 68.69 67.68 73.74
ind_Latn 63.64 52.53 62.63 71.72 27.27 56.57 72.73 69.70 69.70 77.78 73.74 76.77 77.78 57.58
isl_Latn 48.48 37.37 36.36 34.34 25.25 46.46 56.57 53.54 59.60 54.55 66.67 66.67 67.68 55.56
ita_Latn 60.61 54.55 61.62 64.65 35.35 59.60 71.72 66.67 67.68 72.73 76.77 71.72 73.74 70.71
jav_Latn 51.52 58.59 60.61 61.62 31.31 46.46 63.64 63.64 64.65 64.65 69.70 64.65 67.68 73.74
jpn_Jpan 41.41 51.52 64.65 65.66 36.36 56.57 64.65 67.68 66.67 75.76 74.75 69.7 72.73 56.57
kab_Latn 14.14 19.19 17.17 27.27 14.14 31.31 32.32 21.21 26.26 30.30 26.26 25.25 29.29 66.67
kac_Latn 32.32 26.26 34.34 35.35 21.21 34.34 35.35 40.40 39.39 42.42 39.39 41.41 39.39 63.64
kam_Latn 29.29 32.32 32.32 44.44 20.20 29.29 42.42 42.42 39.39 49.49 36.36 40.4 42.42 41.41
kan_Knda 15.15 10.10 59.60 19.19 23.23 45.45 48.48 55.56 56.57 58.59 54.55 55.56 56.57 39.39
kat_Geor 42.42 50.51 20.20 26.26 24.24 45.45 56.57 61.62 59.60 61.62 67.68 62.63 66.67 44.44
kaz_Cyrl 36.36 52.53 35.35 49.49 27.27 47.47 59.60 62.63 62.63 67.68 66.67 64.65 63.64 58.59
kbp_Latn 22.22 25.25 27.27 26.26 21.21 30.30 30.30 38.38 31.31 39.39 37.37 42.42 41.41 51.52
kea_Latn 51.52 39.39 57.58 57.58 29.29 52.53 58.59 60.61 57.58 65.66 64.65 66.67 70.71 64.65
khm_Khmr 31.31 13.13 30.30 29.29 38.38 53.54 58.59 55.56 58.59 65.66 60.61 63.64 67.68 67.68
kik_Latn 28.28 33.33 34.34 31.31 21.21 41.41 39.39 46.46 43.43 50.51 46.46 45.45 50.51 47.47
kin_Latn 31.31 32.32 49.49 30.30 18.18 32.32 44.44 49.49 50.51 52.53 49.49 48.48 54.55 36.36
kir_Cyrl 38.38 50.51 32.32 47.47 29.29 52.53 59.60 61.62 63.64 69.70 60.61 62.63 66.67 71.72
kmb_Latn 28.28 34.34 34.34 43.43 15.15 28.28 41.41 39.39 37.37 39.39 39.39 44.44 42.42 52.53
kmr_Latn 30.30 34.34 38.38 42.42 20.20 45.45 52.53 59.60 57.58 58.59 54.55 54.55 62.63 69.70
kon_Latn 34.34 35.35 44.44 46.46 26.26 43.43 46.46 56.57 51.52 50.51 55.56 48.48 53.54 70.71
kor_Hang 48.48 53.54 44.44 60.61 42.42 62.63 67.68 69.70 68.69 76.77 77.78 69.7 74.75 43.43
lao_Laoo 25.25 10.10 20.20 31.31 31.31 49.49 56.57 61.62 60.61 66.67 63.64 59.6 62.63 58.59
lij_Latn 52.53 51.52 51.52 60.61 23.23 51.52 66.67 64.65 63.64 64.65 68.69 68.69 69.7 54.55
Table 5: Detailed Results (Part I). For previous LLMs, 3-shot results are presented.
Lang LLaMA 2 7B mGPT 13B BLOOM 7B1 XGLM 7.5B MaLA-500
1-shot 2-shot 3-shot 4-shot 5-shot 6-shot 7-shot 8-shot 9-shot 10-shot
lim_Latn 49.49 43.43 49.49 50.51 25.25 47.47 60.61 61.62 60.61 66.67 66.67 64.65 71.72 44.44
lin_Latn 43.43 34.34 41.41 46.46 27.27 43.43 46.46 54.55 56.57 57.58 60.61 54.55 55.56 71.72
lit_Latn 38.38 46.46 39.39 45.45 19.19 50.51 58.59 62.63 65.66 64.65 66.67 63.64 65.66 53.54
lmo_Latn 52.53 48.48 53.54 53.54 19.19 54.55 60.61 54.55 60.61 60.61 60.61 60.61 65.66 63.64
ltz_Latn 56.57 48.48 50.51 50.51 32.32 53.54 65.66 63.64 68.69 70.71 70.71 65.66 72.73 62.63
lua_Latn 35.35 27.27 35.35 39.39 19.19 35.35 41.41 51.52 48.48 47.47 46.46 47.47 48.48 41.41
lug_Latn 28.28 27.27 38.38 36.36 18.18 31.31 41.41 40.40 42.42 44.44 44.44 39.39 44.44 46.46
luo_Latn 31.31 27.27 31.31 36.36 17.17 36.36 32.32 41.41 35.35 37.37 37.37 37.37 44.44 72.73
lus_Latn 42.42 36.36 43.43 44.44 20.20 40.40 43.43 47.47 43.43 47.47 47.47 44.44 49.49 66.67
lvs_Latn 47.47 57.58 44.44 43.43 30.30 59.60 67.68 66.67 64.65 69.70 72.73 68.69 76.77 69.70
mai_Deva 47.47 48.48 62.63 72.73 27.27 51.52 54.55 60.61 54.55 62.63 60.61 54.55 60.61 44.44
mal_Mlym 19.19 39.39 63.64 25.25 17.17 40.40 52.53 51.52 52.53 52.53 49.49 52.53 57.58 56.57
mar_Deva 34.34 47.47 65.66 48.48 21.21 46.46 56.57 60.61 57.58 55.56 60.61 55.56 60.61 72.73
min_Latn 51.52 48.48 59.60 55.56 21.21 52.53 59.60 66.67 56.57 67.68 65.66 58.59 65.66 60.61
mkd_Cyrl 56.57 44.44 39.39 55.56 46.46 57.58 61.62 65.66 68.69 71.72 74.75 70.71 72.73 71.72
mlt_Latn 46.46 43.43 45.45 43.43 20.20 58.59 70.71 68.69 69.70 72.73 74.75 74.75 78.79 68.69
mon_Cyrl 25.25 43.43 28.28 40.40 30.30 48.48 55.56 59.60 61.62 65.66 61.62 60.61 64.65 50.51
mos_Latn 24.24 26.26 30.30 35.35 18.18 29.29 36.36 38.38 35.35 39.39 32.32 37.37 40.4 52.53
mri_Latn 35.35 20.20 27.27 28.28 27.27 39.39 41.41 44.44 40.40 48.48 48.48 43.43 48.48 58.59
mya_Mymr 19.19 49.49 18.18 67.68 18.18 37.37 43.43 41.41 40.40 48.48 48.48 48.48 48.48 43.43
nld_Latn 53.54 47.47 53.54 64.65 38.38 60.61 70.71 74.75 69.70 73.74 77.78 72.73 75.76 75.76
nno_Latn 62.63 49.49 55.56 58.59 26.26 50.51 68.69 62.63 66.67 68.69 69.70 64.65 71.72 66.67
npi_Deva 50.51 41.41 62.63 61.62 28.28 51.52 60.61 65.66 62.63 69.70 66.67 63.64 66.67 67.68
nso_Latn 28.28 29.29 32.32 34.34 18.18 37.37 40.40 42.42 39.39 46.46 46.46 43.43 48.48 60.61
nya_Latn 34.34 33.33 40.40 43.43 22.22 37.37 44.44 53.54 54.55 56.57 56.57 56.57 50.51 50.51
oci_Latn 59.60 49.49 61.62 62.63 26.26 49.49 61.62 61.62 60.61 66.67 65.66 66.67 65.66 66.67
orm_Latn 22.22 23.23 22.22 23.23 14.14 26.26 29.29 35.35 37.37 32.32 38.38 37.37 38.38 58.59
ory_Orya 21.21 11.11 60.61 19.19 19.19 37.37 46.46 49.49 52.53 52.53 57.58 51.52 55.56 54.55
pag_Latn 57.58 47.47 55.56 56.57 29.29 57.58 62.63 69.70 66.67 67.68 67.68 69.7 68.69 57.58
pan_Guru 19.19 11.11 61.62 22.22 21.21 39.39 48.48 54.55 51.52 55.56 55.56 59.6 50.51 73.74
pap_Latn 57.58 49.49 54.55 58.59 25.25 59.60 70.71 70.71 69.70 69.70 72.73 73.74 73.74 71.72
pes_Arab 53.54 44.44 54.55 51.52 30.30 46.46 55.56 58.59 57.58 64.65 64.65 64.65 67.68 68.69
plt_Latn 30.30 30.30 33.33 31.31 20.20 45.45 50.51 52.53 54.55 56.57 60.61 58.59 65.66 74.75
pol_Latn 61.62 48.48 51.52 53.54 36.36 59.60 69.70 66.67 68.69 72.73 71.72 65.66 68.69 71.72
por_Latn 58.59 48.48 58.59 69.70 30.30 58.59 66.67 63.64 66.67 69.70 73.74 67.68 77.78 75.76
prs_Arab 54.55 34.34 48.48 51.52 27.27 49.49 55.56 60.61 60.61 64.65 57.58 61.62 64.65 68.69
pus_Arab 39.39 32.32 41.41 48.48 19.19 38.38 43.43 48.48 48.48 44.44 47.47 47.47 50.51 72.73
quy_Latn 40.40 37.37 43.43 46.46 23.23 43.43 46.46 54.55 47.47 51.52 52.53 48.48 52.53 74.75
ron_Latn 64.65 53.54 60.61 57.58 32.32 57.58 63.64 64.65 67.68 71.72 69.70 63.64 72.73 71.72
run_Latn 29.29 30.30 43.43 33.33 17.17 42.42 52.53 56.57 49.49 55.56 51.52 54.55 54.55 66.67
rus_Cyrl 62.63 51.52 53.54 63.64 43.43 60.61 68.69 67.68 67.68 70.71 73.74 68.69 75.76 74.75
sag_Latn 26.26 29.29 34.34 40.40 24.24 37.37 39.39 44.44 44.44 47.47 43.43 45.45 49.49 44.44
san_Deva 38.38 45.45 52.53 51.52 22.22 41.41 53.54 49.49 42.42 48.48 49.49 42.42 39.39 68.69
scn_Latn 43.43 45.45 52.53 51.52 26.26 60.61 67.68 66.67 66.67 68.69 74.75 70.71 75.76 36.36
sin_Sinh 22.22 09.09 12.12 16.16 23.23 49.49 51.52 53.54 55.56 60.61 59.60 60.61 57.58 63.64
slk_Latn 63.64 52.53 48.48 49.49 32.32 56.57 62.63 62.63 66.67 69.70 68.69 67.68 68.69 63.64
slv_Latn 65.66 53.54 42.42 49.49 32.32 57.58 64.65 65.66 67.68 68.69 70.71 70.71 74.75 63.64
smo_Latn 30.30 30.30 31.31 35.35 24.24 49.49 59.60 60.61 59.60 61.62 65.66 59.6 65.66 76.77
sna_Latn 32.32 26.26 41.41 35.35 21.21 36.36 41.41 43.43 42.42 46.46 47.47 39.39 40.4 68.69
snd_Arab 38.38 35.35 34.34 43.43 25.25 42.42 37.37 45.45 40.40 49.49 47.47 45.45 47.47 55.56
som_Latn 32.32 29.29 26.26 38.38 15.15 31.31 41.41 41.41 47.47 44.44 42.42 45.45 46.46 66.67
sot_Latn 29.29 24.24 35.35 33.33 19.19 40.40 44.44 49.49 50.51 49.49 52.53 51.52 54.55 73.74
spa_Latn 59.60 51.52 66.67 67.68 33.33 58.59 69.70 69.70 67.68 72.73 71.72 73.74 76.77 60.61
srd_Latn 45.45 40.40 49.49 54.55 19.19 54.55 65.66 66.67 62.63 68.69 73.74 66.67 72.73 51.52
srp_Cyrl 62.63 48.48 45.45 59.60 49.49 58.59 64.65 58.59 67.68 69.70 69.70 63.64 69.7 66.67
ssw_Latn 25.25 27.27 33.33 29.29 18.18 36.36 39.39 51.52 49.49 48.48 52.53 47.47 48.48 61.62
sun_Latn 53.54 49.49 57.58 60.61 31.31 59.60 68.69 68.69 67.68 67.68 68.69 66.67 73.74 66.67
swe_Latn 65.66 49.49 52.53 61.62 29.29 57.58 66.67 64.65 66.67 72.73 72.73 73.74 75.76 62.63
swh_Latn 43.43 51.52 60.61 70.71 29.29 53.54 61.62 61.62 58.59 62.63 56.57 62.63 66.67 68.69
szl_Latn 51.52 43.43 44.44 41.41 35.35 52.53 56.57 59.60 64.65 63.64 68.69 62.63 71.72 54.55
tam_Taml 29.29 47.47 59.60 64.65 21.21 39.39 47.47 51.52 55.56 65.66 58.59 57.58 60.61 74.75
tat_Cyrl 36.36 56.57 34.34 47.47 30.30 50.51 55.56 58.59 59.60 61.62 63.64 63.64 67.68 68.69
tel_Telu 22.22 52.53 59.60 59.60 15.15 36.36 48.48 53.54 44.44 54.55 48.48 49.49 51.52 67.68
tgk_Cyrl 26.26 45.45 32.32 45.45 25.25 39.39 49.49 59.60 61.62 61.62 60.61 59.6 63.64 28.28
tgl_Latn 62.63 45.45 45.45 50.51 29.29 61.62 70.71 70.71 67.68 76.77 74.75 71.72 77.78 73.74
tha_Thai 49.49 56.57 23.23 67.68 38.38 56.57 62.63 62.63 68.69 69.70 65.66 66.67 73.74 46.46
tir_Ethi 13.13 14.14 13.13 19.19 17.17 29.29 35.35 29.29 36.36 38.38 38.38 36.36 36.36 72.73
tpi_Latn 62.63 50.51 56.57 55.56 34.34 63.64 73.74 71.72 67.68 75.76 75.76 74.75 76.77 58.59
tsn_Latn 33.33 31.31 36.36 36.36 19.19 37.37 36.36 44.44 40.40 40.40 45.45 45.45 45.45 36.36
tso_Latn 34.34 27.27 29.29 30.30 17.17 33.33 38.38 47.47 43.43 45.45 50.51 45.45 46.46 41.41
tuk_Latn 41.41 43.43 38.38 44.44 25.25 41.41 52.53 54.55 45.45 52.53 53.54 53.54 59.6 68.69
tum_Latn 34.34 31.31 36.36 44.44 24.24 41.41 47.47 53.54 46.46 57.58 51.52 48.48 54.55 46.46
tur_Latn 46.46 50.51 49.49 64.65 21.21 57.58 66.67 66.67 63.64 68.69 69.70 68.69 75.76 74.75
uig_Arab 25.25 17.17 26.26 17.17 19.19 28.28 29.29 27.27 32.32 36.36 37.37 42.42 37.37 57.58
ukr_Cyrl 55.56 52.53 42.42 60.61 39.39 57.58 59.60 60.61 63.64 65.66 64.65 65.66 64.65 54.55
umb_Latn 27.27 29.29 30.30 31.31 17.17 36.36 36.36 36.36 34.34 40.40 38.38 36.36 38.38 44.44
urd_Arab 47.47 42.42 63.64 66.67 24.24 53.54 54.55 55.56 50.51 61.62 60.61 53.54 58.59 56.57
uzb_Latn 39.39 52.53 36.36 37.37 22.22 50.51 56.57 61.62 61.62 68.69 66.67 63.64 70.71 43.43
vec_Latn 52.53 51.52 55.56 63.64 23.23 58.59 65.66 62.63 64.65 69.70 70.71 65.66 73.74 52.53
vie_Latn 62.63 52.53 63.64 68.69 38.38 59.60 65.66 61.62 63.64 69.70 73.74 64.65 73.74 42.42
war_Latn 56.57 49.49 55.56 57.58 34.34 57.58 69.70 73.74 69.70 78.79 75.76 76.77 73.74 44.44
wol_Latn 38.38 34.34 33.33 35.35 21.21 37.37 36.36 41.41 41.41 41.41 43.43 36.36 51.52 64.65
xho_Latn 32.32 31.31 34.34 36.36 24.24 37.37 43.43 47.47 49.49 55.56 51.52 51.52 51.52 65.66
yid_Hebr 18.18 23.23 16.16 18.18 17.17 28.28 34.34 39.39 38.38 38.38 43.43 44.44 41.41 75.76
yor_Latn 21.21 27.27 40.40 34.34 20.20 31.31 37.37 41.41 38.38 43.43 44.44 42.42 41.41 70.71
yue_Hani 41.41 53.54 70.71 67.68 42.42 53.54 68.69 62.63 67.68 69.70 64.65 63.64 68.69 75.76
zsm_Latn 60.61 51.52 60.61 66.67 21.21 58.59 68.69 65.66 71.72 71.72 73.74 72.73 77.78 38.38
zul_Latn 33.33 23.23 35.35 31.31 17.17 36.36 48.48 51.52 49.49 53.54 51.52 52.53 47.47 67.68
all 41.94 39.96 44.22 47.53 26.08 47.05 54.10 55.65 55.34 59.40 59.05 57.26 60.49 60.45
Table 6: Detailed Results (Part II). For previous LLMs, 3-shot results are presented.